OpenStudy (anonymous):

455.9 = volume, find the radius

OpenStudy (accessdenied):

Assuming this is a sphere, we may use the formula for the volume of a sphere and just plug in the volume / solve for the radius.

OpenStudy (accessdenied):

I just want to make sure, this is a sphere, correct?

OpenStudy (accessdenied):

... So, the volume of a sphere is given by the equation: \[ V = \frac{4}{3} \pi r^3 \] Plug in V = 455.9 and solve for r.

OpenStudy (anonymous):

yes it is for a basketball

OpenStudy (anonymous):

thank you so much.

OpenStudy (accessdenied):

You're welcome. :)

OpenStudy (anonymous):

what is the answer?

OpenStudy (accessdenied):

Hmm, did you get an answer yourself?